# Exact Factorization of Unitary Transformations with Spin-Adapted Generators

### v1

@@ Appendix B Seniority 4, intermediate triplet
$A_{63}$ $=-(\hat{n}(i_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})-\hat{n}(i_{\beta})\hat{n}(a_{\alpha})+2\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(a_{\alpha}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{64}$ $=(\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\beta})\hat{n}(b_{\beta})+2\hat{n}(j_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{65}$ $=(\hat{n}(j_{\beta})\hat{n}(b_{\alpha})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\beta})$
$\quad+\hat{n}(j_{\alpha})\hat{n}(b_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\beta})\hat{n}(b_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{66}$ $=(\hat{n}(i_{\beta})\hat{n}(a_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})-\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(a_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(a_{\beta})$
$\quad+\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})\hat{n}(a_{\beta})-\hat{n}(i_{\beta})\hat{n}(a_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{67}$ $=(\hat{n}(a_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(a_{\beta})-\hat{n}(a_{\alpha})\hat{n}(a_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{68}$ $=-(\hat{n}(b_{\beta})-\hat{n}(j_{\beta})\hat{n}(b_{\alpha})-\hat{n}(j_{\beta})\hat{n}(b_{\beta})-\hat{n}(b_{\alpha})\hat{n}(b_{\beta})+2\hat{n}(j_{\beta})\hat{n}(b_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{69}$ $=(\hat{n}(b_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-\hat{n}(b_{\alpha})\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{70}$ $=(\hat{n}(i_{\beta})\hat{n}(j_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(j_{\beta})+\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})\hat{n}(j_{\beta})-\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(j_{\alpha})$
$\quad+\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(j_{\beta})-\hat{n}(i_{\beta})\hat{n}(j_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{71}$ $=(\hat{n}(j_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{72}$ $=(\hat{n}(a_{\beta})\hat{n}(b_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-\hat{n}(a_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\beta})$
$\quad+\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})\hat{n}(b_{\beta})-\hat{n}(a_{\beta})\hat{n}(b_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{73}$ $=(\hat{n}(a_{\beta})-\hat{n}(a_{\beta})\hat{n}(b_{\beta})-\hat{n}(a_{\alpha})\hat{n}(a_{\beta})-\hat{n}(a_{\alpha})\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(a_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{74}$ $=(\hat{n}(i_{\beta})-\hat{n}(i_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\beta})\hat{n}(j_{\beta})-\hat{n}(i_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(j_{\beta}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{75}$ $=(2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(j_{\alpha}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{76}$ $=(\hat{n}(a_{\beta})\hat{n}(i_{\beta})-\hat{n}(a_{\beta})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(i_{\beta})-2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{77}$ $=(\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(j_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{78}$ $=(\hat{n}(i_{\beta})-\hat{n}(a_{\beta})\hat{n}(i_{\beta})-\hat{n}(i_{\beta})\hat{n}(i_{\alpha})+2\hat{n}(a_{\beta})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(a_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{79}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(i_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-2\hat{n}(j_{\alpha})\hat{n}(i_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{80}$ $=(\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})-\hat{n}(a_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{81}$ $=(\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(j_{\beta})\hat{n}(i_{\beta})-\hat{n}(j_{\beta})-2\hat{n}(j_{\beta})\hat{n}(j_{\alpha})\hat{n}(i_{\beta})+\hat{n}(j_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{82}$ $=(\hat{n}(b_{\beta})\hat{n}(b_{\alpha})+\hat{n}(b_{\beta})\hat{n}(a_{\beta})-\hat{n}(b_{\beta})-2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(a_{\beta})+\hat{n}(b_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{83}$ $=(\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(j_{\alpha})-2\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{84}$ $=(\hat{n}(i_{\beta})-\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\beta})+2\hat{n}(b_{\beta})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{85}$ $=(\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})$
@@
$A_{86}$ $=-(\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta})$
@@
$A_{87}$ $=-(\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})-2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\beta})\hat{n}(j_{\alpha}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{88}$ $=(\hat{n}(b_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})-2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(b_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{89}$ $=(\hat{n}(b_{\beta})\hat{n}(b_{\alpha})-\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\beta})+2\hat{n}(b_{\beta})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})$
$A_{90}$ $=(\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\hat{n}(a_{\alpha})-2\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})$
$\quad-\hat{n}(a_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{91}$ $=(2\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})-2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})-\hat{n}(j_{\beta})\hat{n}(j_{\alpha})$
$\quad-\hat{n}(a_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{92}$ $=(2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})+\hat{n}(i_{\beta})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\beta})$
$\quad-\hat{n}(b_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(b_{\alpha}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{93}$ $=(2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{94}$ $=(\hat{n}(a_{\beta})\hat{n}(j_{\beta})-\hat{n}(a_{\beta})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\beta})-2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{95}$ $=(\hat{n}(i_{\beta})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(i_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{96}$ $=-(\hat{n}(a_{\beta})\hat{n}(j_{\beta})-\hat{n}(j_{\beta})+\hat{n}(j_{\beta})\hat{n}(j_{\alpha})-2\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\beta})\hat{n}(j_{\alpha}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
Fourth-Level Commutators:
$A_{97}$ $=(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(a_{\beta})+\hat{n}(j_{\beta})\hat{n}(b_{\alpha})$
@@
$A_{98}$ $=(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})$
@@
$A_{99}$ $=(\hat{n}(a_{\beta})\hat{n}(j_{\beta})-2\hat{n}(b_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})-2\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(j_{\beta})+\hat{n}(j_{\beta})\hat{n}(i_{\beta})$
@@
$A_{100}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})+4\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})$

### v2

@@ Appendix B Seniority 4, intermediate triplet
$A_{57}$ $=-(\hat{n}(i_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(i_{\beta})-\hat{n}(i_{\alpha})\hat{n}(a_{\alpha})-\hat{n}(i_{\beta})\hat{n}(a_{\alpha})+2\hat{n}(i_{\alpha})\hat{n}(i_{\beta})\hat{n}(a_{\alpha}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{58}$ $=(\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(b_{\beta})-\hat{n}(j_{\beta})\hat{n}(b_{\beta})+2\hat{n}(j_{\alpha})\hat{n}(j_{\beta})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{59}$ $=(\hat{n}(b_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})-\hat{n}(a_{\alpha})\hat{n}(b_{\beta})-\hat{n}(b_{\alpha})\hat{n}(b_{\beta})+2\hat{n}(a_{\alpha})\hat{n}(b_{\alpha})\hat{n}(b_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{60}$ $=(\hat{n}(j_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(i_{\alpha})\hat{n}(j_{\beta})-\hat{n}(j_{\alpha})\hat{n}(j_{\beta})+2\hat{n}(i_{\alpha})\hat{n}(j_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{61}$ $=(2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(j_{\alpha}))\hat{G}_{i_{\alpha}a_{\beta}}^{i_{\beta}a_{\alpha}}$
$A_{62}$ $=(\hat{n}(a_{\beta})\hat{n}(i_{\beta})-\hat{n}(a_{\beta})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(i_{\beta})-2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{j_{\alpha}b_{\beta}}^{j_{\beta}b_{\alpha}}$
$A_{63}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(i_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(i_{\beta})-2\hat{n}(j_{\alpha})\hat{n}(i_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{a_{\alpha}b_{\beta}}^{a_{\beta}b_{\alpha}}$
$A_{64}$ $=(\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})-\hat{n}(a_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(a_{\beta}))\hat{G}_{i_{\alpha}j_{\beta}}^{i_{\beta}j_{\alpha}}$
$A_{65}$ $=(\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(j_{\alpha})-2\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{66}$ $=(\hat{n}(i_{\beta})-\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\beta})+2\hat{n}(b_{\beta})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{67}$ $=(\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})-\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(i_{\beta})\hat{n}(i_{\alpha})$
@@
$A_{68}$ $=-(\hat{n}(a_{\alpha})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\beta})\hat{n}(j_{\beta})\hat{n}(j_{\alpha})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta})$
@@
$A_{69}$ $=-(\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})-2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\beta})\hat{n}(j_{\alpha}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{70}$ $=(\hat{n}(b_{\beta})\hat{n}(i_{\beta})-\hat{n}(b_{\beta})-2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\beta})+\hat{n}(b_{\beta})\hat{n}(b_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\beta}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{71}$ $=(2\hat{n}(b_{\beta})\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})-\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(b_{\alpha})-\hat{n}(b_{\beta})\hat{n}(i_{\alpha}))\hat{G}_{j_{\alpha}a_{\beta}}^{j_{\beta}a_{\alpha}}$
$A_{72}$ $=(\hat{n}(a_{\beta})\hat{n}(j_{\beta})-\hat{n}(a_{\beta})+\hat{n}(a_{\beta})\hat{n}(a_{\alpha})+\hat{n}(a_{\alpha})\hat{n}(j_{\beta})-2\hat{n}(a_{\beta})\hat{n}(a_{\alpha})\hat{n}(j_{\beta}))\hat{G}_{i_{\alpha}b_{\beta}}^{i_{\beta}b_{\alpha}}$
$A_{73}$ $=(1-\hat{n}(i_{\alpha})-\hat{n}(j_{\beta})-\hat{n}(a_{\beta})-\hat{n}(b_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(b_{\alpha})+\hat{n}(i_{\alpha})\hat{n}(a_{\beta})+\hat{n}(j_{\beta})\hat{n}(b_{\alpha})$
@@
$A_{74}$ $=(1-\hat{n}(i_{\beta})-\hat{n}(j_{\alpha})-\hat{n}(a_{\alpha})-\hat{n}(b_{\beta})+\hat{n}(j_{\alpha})\hat{n}(a_{\alpha})+\hat{n}(j_{\alpha})\hat{n}(b_{\beta})+\hat{n}(i_{\beta})\hat{n}(a_{\alpha})$
@@
$A_{75}$ $=(\hat{n}(a_{\beta})\hat{n}(j_{\beta})-2\hat{n}(b_{\beta})\hat{n}(j_{\beta})\hat{n}(i_{\beta})-2\hat{n}(b_{\beta})\hat{n}(a_{\beta})\hat{n}(j_{\beta})+\hat{n}(j_{\beta})\hat{n}(i_{\beta})$
@@
$A_{76}$ $=(\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})+\hat{n}(b_{\alpha})\hat{n}(i_{\alpha})+4\hat{n}(b_{\alpha})\hat{n}(a_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})-2\hat{n}(b_{\alpha})\hat{n}(j_{\alpha})\hat{n}(i_{\alpha})$